If the demand fun

tion is

P = Q

2 − 7Q − 8,

where P and Q are the pri

e and quantity respe

tively, determine the point pri

e elasti

ity of demand

(rounded to three de

imal pla

es) if P = 51. Is demand elasti

 or inelasti

 at this pri

e?

[1℄ ed = −0.075; inelasti


[2℄ ed = −0.530; elasti


[3℄ ed = 1.300; elasti


[4℄ ed = −0.255; inelasti


1
Expert's answer
2021-06-29T07:11:02-0400

f the demand function is P = Q2 - 7Q - 8, where P and Q are the price and quantity respectively, determine the point priсe elasticity of demand (rounded to three decimal places) if P = 51. Is demand elastic or inelastic at this price?


"\frac{\varDelta p}{\varDelta Q}=2Q-7"

"E=\frac{\varDelta Q}{\varDelta p}\times \frac{p}{Q}"


"E=0.25"

4℄ ed = −0.255; inelasti